OpenStudy (anonymous):

Help please 3/12 Divided by 5/6

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

I can help Do you know how to divide fractions

OpenStudy (anonymous):

Kinda

OpenStudy (anonymous):

first simplify 3/12

OpenStudy (anonymous):

what did you get

OpenStudy (dakid88):

first get a common denominator

OpenStudy (misty1212):

\[\frac{3}{12}\div \frac{5}{6}=\frac{3}{12}\times \frac{6}{5}\] cancel first, multiply last some elementary school math teachers call this KFC Keep the first one Flip the second Change the \(\div\) to a \(\times\)

OpenStudy (anonymous):

dakid88 you dont need a common denominator

OpenStudy (dakid88):

i mean first simplify 3/12

OpenStudy (dakid88):

3/12 would be simplified to 1/4

OpenStudy (dakid88):

next simplify 5/6

OpenStudy (dakid88):

and tht would simplify to 5/6

OpenStudy (dakid88):

then multiply those and u get wht @hunter123123

OpenStudy (mathstudent55):